|2x -4| = 12


thanks for all the help!

Answer by joecbaseball(37)

When dealing with absolute values, you need to know that what comes out of the absval sign is always positive. Hence, you need to consider that. This is what I mean:
The right hand side is = 12, so the left hand side must = 12 OR -12
So, set the left hand side = to both 12 and -12.
2x-4 = 12, and solve for x
Add 4 to both sides to get 2x = 16, and divide both sides by 2 to get x by itself:
x=16/2 = 8. So, 8 is one of your answers.
Now, set the left hand side = -12:
2x-4 = -12
Add 4 to both sides to get 2x = -8, and divide both sides by 2 to get x by itself:
x=-8/2 = -4. So -4 is the other answer.
So, this problem has two solutions, 8 and -4.
Good luck,
JoeC